# Exploring Performance-Productivity Trade-offs in AMT Runtimes: A Task Bench Study of Itoyori, ItoyoriFBC, HPX, and MPI

### v1

@@ 6 Conclusions
We presented the first Task Bench evaluation of the Itoyori and ItoyoriFBC runtime systems. Through comparison against MPI and HPX across diverse scenarios, we identified distinct performance and productivity characteristics.
Itoyori demonstrates that the PGAS model combined with global fork-join parallelism can achieve high efficiency in communication-intensive workloads, outperforming MPI in all-to-all patterns. Its global address space also significantly improves programmer productivity, reducing code size by nearly 50% compared to MPI. However, its random work stealing scheduler incurs overheads at low task granularities, limiting effectiveness for fine-grained, regular workloads.
ItoyoriFBC introduces a flexible future-based model that supports arbitrary task graphs. While its current implementation incurs overheads from explicit task spawning and future management, it offers a promising direction for irregular applications that cannot be easily expressed with nested fork-join structures.
Moreover, we improved the HPX Task Bench implementation. Results indicate that HPX excels at intra-node load balancing but faces scalability challenges in communication-heavy scenarios.
Future research should investigate an improved ItoyoriFBC implementation that does not require barriers or future re-wrapping.
@@ Acknowledgement
This research was partially funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under project numbers 12345 and 6789.

### v2

@@ 6 Conclusions
We presented the first Task Bench implementations of the Itoyori and ItoyoriFBC runtime systems and compared them with MPI and HPX across diverse scenarios, identifying distinct performance and productivity trade-offs.
The Itoyori implementation demonstrates that the PGAS model combined with global fork-join parallelism can achieve high efficiency in communication-intensive workloads, outperforming the MPI implementation in all-to-all patterns. Its global address space also significantly improves programmer productivity, reducing code size by nearly 50% compared to the MPI implementation. However, its random work stealing scheduler incurs overheads at low task granularities, limiting effectiveness for fine-grained, regular workloads.
The ItoyoriFBC implementation introduces a flexible future-based model supporting arbitrary task graphs. While its current implementation incurs overheads from explicit task spawning and future management, it offers a promising direction for irregular applications not easily expressed with NFJ structures.
Moreover, we improved the HPX Task Bench implementation. Results indicate that the HPX implementation excels at intra-node load balancing but faces scalability challenges in communication-heavy scenarios.
Future research should investigate an improved ItoyoriFBC implementation that does not require barriers or future re-wrapping. Additionally, alternative HPX implementations using nested fork-join or future-based paradigms could help isolate the impact of programming style from runtime characteristics. Strong scaling analysis also remains relevant future work. In this work, we used AMT’s default work-stealing algorithm and parameter settings. A sensitivity study of these parameters is needed to analyze their influence on performance.
@@ Acknowledgement
This research was partially funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under project numbers 512078735 and 558599020.